Given the demand for a product as Qd = 100 - 5P and the supply is given as Qs = -20 + 3p. You are told equilibrium is obtained at the point where Qd = Qs. The equilibrium price for the product is

1
Expert's answer
2021-06-13T17:29:01-0400

"Qd=Qs\\100-5p=-20+3p\\100+20=3p+5p\\120=8p\\p=15"


The equilibrium price for the product is 15